If the shortest wave length of Lyman series of H atom is x, then the wave length of Balmer series of H atom will be 1) 9x/52) 36x/53) 5x/94) 5x/36Explanation please..Thank You

Grade:11

1 Answers

Vikas TU
7 years ago
If the shortest wave length of Lyman series of H atom is x then,
1/k = R*1^2(1/1^2 – 1/2^2)
x = 4/3R...................(1)
For Balmer series,
1/y = R*1^2(1/4 – 1/9)
y = 36/5R.......................(2)
dividing eqns. (1) and (2) we get,
x/y = 20/36 = > 10/18 => 5/9
y = 9x/5